Hello, thanks in advanced:
The sum of the perimeters of two different squares is 32 centimeters, and the difference between their perimeters is 8 centimeters. If x represents the side length of the larger square and y represents the side length of the smaller square,write 2 equations representing this.

Answer by rothauserc(4718)

4x + 4y = 32
4x - 4y = 8
************************
we can reduce the above two equations to
x + y = 8
x - y = 2
*************************
we can add the above two equations and determine the value of x
2x = 10
x = 5, therefore y = 3